# CLF-RL: Control Lyapunov Function Guided Reinforcement Learning

### v1

@@ I Introduction
Traditional model-based control frameworks address these challenges through principled formulations such as trajectory optimization. Methods based on hybrid zero dynamics (HZD) [1] and feedback-linearization have enabled offline generation of stable walking gaits under physical and task constraints [2]. These techniques provide strong guarantees on stability and constraint satisfaction but require solving large-scale nonlinear programs that are computationally demanding and sensitive to model mismatch, limiting their practicality for deployment. To enable faster and more reactive planning, many approaches leverage reduced order models such as the linear inverted pendulum (LIP) and its variants [3, 4], which simplify the dynamics by focusing on the center of mass and foot placement. While computationally efficient and suitable for stabilization and mid-level planning, these models rely on strong assumptions such as massless legs and constant center of mass height, which limit their ability to generate highly dynamic and agile whole body motion.
@@
Recent advances in computation and GPU-parallelizable physics simulators have made reinforcement learning (RL) a promising alternative for locomotion control. By offloading the heavy computation to offline training in massively parallel simulations, RL enables lightweight policy inference during online deployment to hardware [5, 6, 7, 8, 9]. However, applying RL to bipedal locomotion remains challenging due to its reliance on heuristic reward design, which is tedious to construct, sensitive to tuning, and, if poorly shaped, can lead to unstable gaits, prolonged training, and poor sim-to-real transfer. A strategy to alleviate the burden of this process is to embed model-based reference trajectories directly into the RL reward. This idea has been explored using both reduced order models [10, 11] and full-order models within the HZD framework [12]. [13] proposes a different way to incorporate a LIP controller into RL by generating desired footstep locations during training and using it to provide feedback control for rough terrain.
@@
While prior model-guided approaches improve structure during training, they typically reward only the instantaneous tracking error, treating two states with the same error magnitude equally—even if one is actively converging toward the reference and the other is diverging. This can cause policies to undervalue transient corrections. To address these limitations, we propose a reward shaping framework built on control Lyapunov functions (CLFs), a fundamental tool in nonlinear control theory for generating certifiably stable controllers [14, 15, 16]. CLFs have previously been used in bipedal locomotion, often in conjunction with HZD, to formally guarantee the stability of periodic gaits [17] and have also been integrated into learning-based methods in other domains [18, 19]. For instance, [20] introduces a reward-reshaping method that incorporates a candidate CLF for fine-tuning policies using minimal hardware data.
@@
Our approach embeds both the CLF and its associated decrease condition directly into the RL reward, where the decrease condition encodes a desired minimal convergence rate, providing meaningful intermediate rewards whenever the error is being reduced, similar to potential-based reward shaping, which has been shown to be more robust to scaling and hence simplify tuning [21]. This yields a reward component that is principled, easy to integrate, and less sensitive to manual reward balancing than conventional tracking formulations.
@@
Building on these properties, we propose a structured reward shaping approach that combines model-based reference planning with CLF-inspired objectives to guide policy learning toward stable and robust behaviors (Fig. 1). We use two complementary planners for generating velocity-conditioned reference trajectories: (1) a reduced-order linear inverted pendulum (LIP) model for online generation, and (2) a precomputed hybrid zero dynamics (HZD) gait library from full-body offline trajectory optimization. Both produce periodic orbits but differ in fidelity and real-time suitability. From these trajectories, we construct a CLF $V=\eta^{\top}P\eta$, where $\eta$ denotes the output tracking error. Once a reference trajectory is available, implementing the CLF decrease condition requires minimal effort: the tuning process is straightforward, involving far fewer reward terms than conventional rewards and resembling gain tuning in a tracking controller. The CLF and its associated stability condition is embedded into the RL reward to shape the policy to promote stable behaviors during training. We validate our approach through both simulation and hardware experiments on the Unitree G1 humanoid robot. Results demonstrate that CLF-based reward shaping improves tracking performance at high velocities, reduces variance under randomized model perturbations, and enhances robustness compared to standard RL baselines—offering a promising path toward more robust and theoretically grounded locomotion learning.
@@
Figure 2: Overview of the proposed CLF-guided reinforcement learning framework. A desired velocity $v^{d}$ is passed to a reference generator (e.g., H-LIP or HZD) to produce targets $y^{d}_{\alpha},\dot{y}^{d}_{\alpha}$. These, along with the robot state and privileged variables $o_{t}^{\text{priv}}$, are used to compute a CLF-based reward. The actor-critic policy is trained with this reward and outputs joint targets $q_{\text{target}}$ for the robot.
@@ II Preliminaries
Consider a robotic system with generalized coordinates $q=[q_{b}^{\top},\,q_{a}^{\top}]^{\top}\in\mathcal{Q}\subset\mathbb{R}^{n}$, where $q_{b}\in SE(3)$ denotes the floating-base coordinates and $q_{a}\in\mathbb{R}^{m}$ the actuated degrees of freedom (DoFs). The control input is $u\in\mathbb{R}^{m}$, and the full-order state is given by $x=[q^{\top},\,\dot{q}^{\top}]^{\top}$.
II-A Hybrid Dynamics
Legged locomotion is often modeled as a hybrid system with continuous dynamics and discrete transitions such as impact events across domains [1]. Let $\mathcal{D}$ denote the domain and $\mathcal{S}$ the guard of the system. The hybrid system $\mathcal{H}$ can be described as:
@@ II-A Hybrid Dynamics
II-B Periodic Gait Generation
Legged locomotion is inherently periodic, alternating between stance and swing phases. This motivates analysis via step-to-step (S2S) dynamics, expressed by the discrete-time Poincaré return map $P:\mathcal{S}\to\mathcal{S}$, where $x_{k+1}=P(x_{k})$ evaluates the state at successive impacts. A periodic gait corresponds to a fixed point $x^{*}$ such that $P(x^{*})=x^{*}$ and is exponentially stable if this fixed point is locally attracting [22].
Periodic gaits can be generated by shaping the return map to admit a stable fixed point. Here, we focus on two complementary approaches: full-order hybrid zero dynamics (HZD) and H-LIP models.
Hybrid Zero Dynamics: The Hybrid Zero Dynamics (HZD) framework is a model-based approach for designing stable walking gaits for bipedal robots. It operates by enforcing a set of virtual constraints, typically specified using Bézier polynomials with coefficients $\alpha$. These constraints define desired trajectories for joint or end-effector motion as functions of a monotonically increasing phase variable. Driving the output of the virtual constraints to zero renders the zero dynamics manifold, denoted $\mathcal{Z}^{\alpha}$, invariant and attractive.
To achieve periodic walking under hybrid dynamics, the manifold must also be impact invariant, meaning that the system returns to $\mathcal{Z}^{\alpha}$ after a discrete event (e.g., foot impact):
@@ II-B Periodic Gait Generation
Hybrid Linear Inverted Pendulum: Assuming constant angular momentum and a fixed $\textrm{p}_{\textrm{com}}^{z}$ leads to the well-known Linear Inverted Pendulum (LIP) model, widely used in humanoid walking control [3, 23].
To account for the hybrid nature of bipedal walking, [24] proposed the Hybrid Linear Inverted Pendulum (H-LIP) model. Like the LIP, it assumes constant CoM height and stepping frequency, but it explicitly models the hybrid walking dynamics of underactuated bipeds. The deviation between the full-order model and H-LIP is treated as a bounded error, enabling the design of state-feedback stepping controllers that stabilize horizontal CoM states within bounded invariant sets. The H-LIP dynamics for the single support phase (SSP) and double support phase (DSP) are given by:
$\mathnormal{a}_{\textrm{com}}^{\{x,y\}}=\frac{g}{\textrm{p}_{\textrm{com}}^{z}}\textrm{p}_{\textrm{com}}^{\{x,y\}}\text{ (SSP)}\quad\quad\quad\mathnormal{a}_{\textrm{com}}=0\text{ (DSP)}$
where the $\textrm{p}_{\textrm{com}}^{z}$ and domain duration ($T_{\textrm{SSP}}$, $T_{\textrm{DSP}}$) are constant. Assuming no velocity jump, we can combine the two domains with step size $u^{\textrm{H-LIP}}$ equivalently as:
@@
H-LIP model S2S dynamics are given by $x_{k+1}^{\textrm{H-LIP}}=Ax_{k}^{\textrm{H-LIP}}+Bu^{\textrm{H-LIP}}_{k}$, with $x^{\textrm{H-LIP}}=[\textrm{p}_{\textrm{com}},\textrm{v}_{\textrm{com}}]^{\top}$, which can be used to approximate the Poincaré return map of the full order system. For more details on the H-LIP model formulation and derivation of its step-to-step dynamics, we refer the reader to [24].
Remarks: While H-LIP enables fast online gait generation with foot placement feedback, our method uses only nominal planned trajectories, obtained by identifying the fixed point and computing the continuous trajectory analytically. Although HZD operates on full order dynamics and H-LIP relies on reduced order approximations, both regulate the return map to achieve stable periodic gaits, a shared focus that motivates our reward shaping framework.
II-C Control Lyapunov Functions (CLFs)
For a nonlinear control-affine system of the form $\dot{x}=f(x)+g(x)u$, a continuously differentiable, positive definite function $V(x):\mathbb{R}^{2n}\rightarrow\mathbb{R}$ is called an exponentially stabilizing control Lyapunov function (CLF) if there exists a control input $u$ such that
@@ II-C Control Lyapunov Functions (CLFs)
for scalar $\lambda>0$. This inequality guarantees that $V(x)$ decreases along system trajectories, implying exponential convergence of the state $x$ to the origin (or a desired manifold).
For hybrid systems such as bipedal locomotion with discrete impact events, CLFs can still be applied within each continuous domain of the dynamics. Prior work has shown that if the rate of convergence in the continuous phase is sufficiently fast, it can offset the destabilizing effects of impacts [17], providing a principled way to certify stability despite hybrid transitions.
Importantly, the CLF condition (6) defines a set of admissible control inputs rather than a unique control law. Any control input $u$ that satisfy the inequality contributes to stability. This property makes CLFs especially appealing for reinforcement learning: rather than enforcing a fixed feedback controller, we can embed the CLF condition directly into the reward function, encouraging the policy to learn Lyapunov-stable behavior, without constraining it to a specific control structure.
@@ III Reference-Guided Reward Shaping
We consider the reinforcement learning (RL) problem of learning a locomotion policy $\pi_{\theta}(a_{t}|o_{t})$ that maps proprioceptive observations $o_{t}$ to actions $a_{t}$. The policy is trained using Proximal Policy Optimization (PPO) [25], a widely used actor-critic method that optimizes a clipped surrogate objective with entropy regularization. Instead of relying solely on sparse or handcrafted rewards, we embed model-based prior knowledge into the reward through a control Lyapunov function (CLF), promoting stability with respect to a reference trajectory. The overall framework, including reference generation and CLF-based reward shaping, is illustrated in Fig. 2.
@@ III-A Reference Trajectory Generation
Our approach assumes access to a nominal reference trajectory that encodes a periodic or quasi-periodic walking motion given a desired velocity. In this work, we focus on two sources of reference trajectories: the H-LIP model and full-order HZD optimization. However, the framework is agnostic to the origin of the trajectory.
H-LIP-Based Reference: Given a target walking velocity $v^{d}$, we first solve for the H-LIP fixed point $x^{*}=Ax^{*}+Bu^{*}$, where the step size $u^{*}=v^{d}T$ corresponds to the desired walking speed and nominal step time. This yields a periodic center-of-mass (CoM) state $x^{*}=[p_{\text{com}},v_{\text{com}}]^{\top}$ consistent with the H-LIP step-to-step dynamics. Using this fixed point, we analytically generate the CoM trajectory over one step. The swing foot trajectory is constructed using a time-parameterized 5th-order Bézier curve to ensure foot clearance and smooth touchdown. For other end-effector or joint trajectories (e.g., torso pitch, knees, or arms), we use simple sinusoidal patterns that are synchronized with the gait phase. This procedure produces time-varying desired trajectories for the end effectors and joints.
@@
$\alpha^{*}$ $=\mathop{\arg\min}_{\alpha}\;J(\alpha)$
s.t. $\text{Closed-loop dynamics},$

### v2

@@ I Introduction
Traditional model-based control frameworks address these challenges through principled formulations such as trajectory optimization. Methods based on hybrid zero dynamics (HZD) [westervelt2003hybrid] and feedback-linearization have enabled offline generation of stable walking gaits under physical and task constraints [reher2020algorithmic]. These techniques provide strong guarantees on stability and constraint satisfaction but require solving large-scale nonlinear programs that are computationally demanding and sensitive to model mismatch, limiting their practicality for deployment. To enable faster and more reactive planning, many approaches leverage reduced-order models such as the linear inverted pendulum (LIP) and its variants [kajita2003biped, kajita_3d_2001], which simplify the dynamics by focusing on the center of mass and foot placement. While computationally efficient and suitable for stabilization and mid-level planning, these models rely on strong assumptions such as massless legs and constant center of mass height, which limit their ability to generate highly dynamic and agile whole body motion.
@@
Recent advances in computation and GPU-parallelizable physics simulators have made reinforcement learning (RL) a promising alternative for locomotion control. By offloading the heavy computation to offline training in massively parallel simulations, RL enables lightweight policy inference during online deployment to hardware [lee2020learning, xie2020learning, rudin_learning_2022, zhuang_humanoid_2024, li_reinforcement_2024]. However, applying RL to bipedal locomotion remains challenging due to its reliance on heuristic reward design, which is tedious to construct, sensitive to tuning, and, if poorly shaped, can lead to unstable gaits, prolonged training, and poor sim-to-real transfer. A strategy to alleviate the burden of this process is to embed model-based reference trajectories directly into the RL reward. This idea has been explored using reduced-order models [batke_optimizing_2022, green_learning_2021], full-order models [liu_opt2skill_2025], and with the HZD framework [li_reinforcement_2021]. [lee_integrating_2024] proposes a different way to incorporate a LIP controller into RL by generating desired footstep locations during training and using it to provide feedback control for rough terrain.
@@
While prior model-guided approaches improve structure during training, they typically reward only the instantaneous tracking error, treating two states with the same error magnitude equally—even if one is actively converging toward the reference and the other is diverging. This can cause policies to undervalue transient corrections. To address these limitations, we propose a reward shaping framework built on control Lyapunov functions (CLFs), a fundamental tool in nonlinear control theory for generating certifiably stable controllers [sontag_lyapunov-like_1983, sontag_universal_nodate, artstein_stabilization_1983]. CLFs have previously been used in bipedal locomotion, often in conjunction with HZD, to formally guarantee the stability of periodic gaits [ames_rapidly_2014] and have also been integrated into learning-based methods in other domains [westenbroek_combining_2021, choi_reinforcement_2020]. For instance, [westenbroek_lyapunov_2022] introduces a reward-reshaping method that incorporates a candidate CLF for fine-tuning policies using minimal hardware data.
@@
Our approach embeds both the CLF and its associated decrease condition directly into the RL reward, where the decrease condition encodes a desired minimal convergence rate, providing meaningful intermediate rewards. Although the idea may seem similar to potential shaping [jeon2023benchmarking], there are fundamental differences. Potential shaping replaces the original reward, while we propose an additional reward, and this addition is designed directly according to the structure of a valid CLF. Further, our proposed framework reduces all the tracking to a single Lyapunov function, which as we discuss later, experimentally, does not work well with potential shaping. Overall, the proposed reward formulation is principled, easy to integrate, and less sensitive to manual reward balancing than conventional tracking formulations.
@@
Building on these properties, we propose a structured reward shaping approach that combines model-based reference planning with CLF-inspired objectives to guide policy learning toward stable and robust behaviors (Fig. 1). We use two complementary planners for generating velocity-conditioned reference trajectories: (1) a reduced-order linear inverted pendulum (LIP) model for online generation, and (2) a precomputed hybrid zero dynamics (HZD) gait library from full-body offline trajectory optimization. Both produce periodic orbits but differ in fidelity and real-time suitability. From these trajectories, we construct a CLF $V=\eta^{\top}P\eta$, where $\eta$ denotes the output tracking error. Once a reference trajectory is available, implementing the CLF decrease condition requires minimal effort: the tuning process is straightforward, involving far fewer reward terms than conventional training pipelines and resembling gain tuning in a tracking controller. The CLF and its associated stability condition is embedded into the RL reward to shape the policy to promote stable behaviors during training. We validate our approach through both simulation and hardware experiments on the Unitree G1 humanoid robot. Results demonstrate that CLF-based reward shaping improves tracking performance, reduces variance under randomized model perturbations, and enhances robustness compared to standard RL baselines—offering a promising path toward more robust and theoretically grounded locomotion learning.
@@
Figure 2: Overview of CLF-guided reinforcement learning framework. A desired velocity $v^{d}$ is passed to a reference generator (e.g., H-LIP or HZD) to produce targets $y^{d}_{\alpha},\dot{y}^{d}_{\alpha}$. These, along with the robot state and privileged variables $o_{t}^{\text{priv}}$, are used to compute a CLF-based reward. The actor-critic policy is trained with this reward and outputs joint targets $q_{\text{target}}$ for the robot.
@@ II Preliminaries
Consider a robotic system with generalized coordinates $q=[q_{b}^{\top},\,q_{a}^{\top}]^{\top}\in\mathcal{Q}\subset\mathbb{R}^{n}$, where $q_{b}\in SE(3)$ denotes the floating-base coordinates and $q_{a}\in\mathbb{R}^{m}$ the actuated degrees of freedom (DoFs). The control input is $u\in\mathbb{R}^{m}$, and the full-order state is given by $x=[q^{\top},\,\dot{q}^{\top}]^{\top}$. Legged locomotion is often modeled as a hybrid system with continuous dynamics and discrete transitions such as impact events across domains [westervelt2003hybrid]. Let $\mathcal{D}$ denote the domain and $\mathcal{S}$ the guard of the system. The hybrid system $\mathcal{H}$ can be described as:
@@
II-A Periodic Gait Generation
Legged locomotion is fundamentally periodic, alternating between stance and swing phases. This natually motivates a step-to-step (S2S) analysis via the discrete-time Poincaré return map $P:\mathcal{S}\to\mathcal{S}$, with $x_{k+1}=P(x_{k})$ giving the state from one impact to the next. A periodic gait is a fixed point $x^{*}$ such that $P(x^{*})=x^{*}$ and is exponentially stable when this fixed point is locally attracting [morris2005restricted]. Periodic motions can thus be synthesized by shaping the return map to yield a stable fixed point. Here, we focus on two complementary approaches: hybrid zero dynamics (HZD) and H-LIP models.
Hybrid Zero Dynamics: The Hybrid Zero Dynamics (HZD) approach synthesizes stable bipedal gaits by enforcing a set of phase-based virtual constraints, typically specified using Bézier polynomials with coefficients $\alpha$. These constraints define desired trajectories for joint or end-effector motion as functions of a monotonically increasing phase variable. Driving the output of the virtual constraints to zero renders the zero dynamics manifold, denoted $\mathcal{Z}^{\alpha}$, invariant and attractive. To achieve periodic walking under hybrid dynamics, the manifold must also be impact invariant, meaning that the system returns to $\mathcal{Z}^{\alpha}$ after a discrete event (e.g., foot impact):
@@ II-A Periodic Gait Generation
Hybrid Linear Inverted Pendulum: Assuming constant (centroidal) angular momentum and fixed $\textrm{p}_{\textrm{com}}^{z}$ yields the well-known Linear Inverted Pendulum (LIP) for humanoid walking [kajita2003biped, kajita2002realtime]. To capture the hybrid nature of the gait, [xiong20223] introduced the Hybrid LIP (H-LIP), which, like LIP, assumes constant CoM height and step timing, but explicitly models single (SSP) and double-support (DSP) domains with a reset map. The mismatch to the full-order dynamics is treated as a bounded disturbance, enabling state-feedback foot-placement laws that render the horizontal CoM states bounded and invariant. Assuming no velocity jump, the two domains with step size $u^{\textrm{H-LIP}}$ compose into the equivalent reset:
@@
The S2S dynamics then take the linear form $x_{k+1}^{\textrm{H-LIP}}=Ax_{k}^{\textrm{H-LIP}}+Bu^{\textrm{H-LIP}}_{k}$, with $x^{\textrm{H-LIP}}=[\textrm{p}_{\textrm{com}},\textrm{v}_{\textrm{com}}]^{\top}$, which serves as a reduced-order approximation of the Poincaré map of the full-order system; see [xiong20223] for derivation details.
Remarks: Although H-LIP enables online, feedback-based foot placement, we use steady-state trajectories derived by fixed-point identification and analytic rollout. Regardless of model order (full-order HZD vs. reduced-order H-LIP), both methods stabilize periodic gaits by regulating the Poincaré map, motivating our reward-shaping framework.
II-B Control Lyapunov Functions (CLFs)
For a nonlinear control-affine system of the form $\dot{x}=f(x)+g(x)u$, a continuously differentiable, positive definite function $V(x):\mathbb{R}^{2n}\rightarrow\mathbb{R}$ is called an exponentially stabilizing control Lyapunov function (CLF) if there exists a control input $u$ such that, for some scalar $\lambda>0$,
@@ II-B Control Lyapunov Functions (CLFs)
This inequality guarantees that $V(x)$ decreases along system trajectories, implying exponential convergence of the state $x$ to the origin (or a desired manifold). Importantly, the CLF condition (6) defines a set of admissible control inputs rather than a unique control law. Any control input $u$ that satisfies the inequality contributes to stability. This property makes CLFs especially appealing for reinforcement learning: rather than enforcing a fixed feedback controller, we can embed the CLF condition directly into the reward function, encouraging the policy to learn Lyapunov-stable behavior, without constraining it to a specific control action. For hybrid systems such as bipedal locomotion with discrete impact events, CLFs can be applied only within each continuous domain of the dynamics. Prior work has shown that if the rate of convergence in the continuous phase is sufficiently fast, it can offset the destabilizing effects of impacts [ames_rapidly_2014], providing certifiable stability.
@@ III Reference-Guided Reward Shaping
We consider the reinforcement learning (RL) problem of learning a locomotion policy $\pi_{\theta}(a_{t}|o_{t})$ that maps proprioceptive observations $o_{t}$ to actions $a_{t}$ using Proximal Policy Optimization (PPO) [schulman2017proximal]. Instead of relying solely on handcrafted rewards, we embed model-based prior knowledge into the reward through a control Lyapunov function (CLF), promoting stability with respect to a reference trajectory. The full framework, including reference generation and CLF-based reward shaping, is illustrated in Fig. 2.
@@ III-A Reference Trajectory Generation
Our approach assumes a nominal reference trajectory that encodes a periodic or quasi-periodic walking motion given a desired velocity. We focus on two sources of reference trajectories: the H-LIP model and full-order HZD optimization, but the framework is agnostic to the origin of the trajectory.
H-LIP-Based Reference: Given a target walking velocity $v^{d}$, we first solve for the H-LIP fixed point $x^{*}=Ax^{*}+Bu^{*}$, where the step size $u^{*}=v^{d}T$ corresponds to the desired walking speed and nominal step time. This yields a periodic center-of-mass (CoM) state $x^{*}=[p_{\text{com}},v_{\text{com}}]^{\top}$ consistent with the H-LIP S2S dynamics. Using this fixed point, we analytically generate the CoM trajectory over one step. The swing foot trajectory is constructed using a time-parameterized 5th-order Bézier curve to ensure foot clearance and smooth touchdown. For other end-effector or joint trajectories, we use simple sinusoidal patterns that are synchronized with the gait phase to produce time-varying desired trajectories.
@@
$\{\alpha^{*},X^{*}\}=$ $\operatornamewithlimits{argmin}_{\alpha,X}\Phi(X)$
s.t. $\dot{x}=f(x)+g(x)u$ (Dynamics)
$\Delta(\mathcal{S}\cap\mathcal{Z}_{\alpha})\subset\mathcal{Z}_{\alpha}$ (HZD Condition)
$X_{\text{min}}\leq X\leq X_{\text{max}}$ (Decision Variables)
$c_{\text{min}}\leq c(X)\leq c_{\text{max}},$ (Physical Constraints)
where $X=(x_{0},u_{0}...,x_{N},u_{N})\in\mathbb{R}^{n_{d}}$ denotes the collection of decision variables, with $n_{d}\in\mathbb{N}$. Each $x_{i}$ represents the state at node $i$. The cost function is given $\Phi:\mathbb{R}^{n_{d}}\to\mathbb{R}_{\geq 0}$, while the constraints are represented by $c:\mathbb{R}^{n_{d}}\to\mathbb{R}^{n_{p}}$ with $n_{p}\in\mathbb{N}$. These constraints encode the physical laws governing locomotion, such as friction cone conditions, workspace limits, and actuator capacity [reher2020algorithmic]. Solving this optimization yields a stable periodic solution to the walking dynamics for a single step, parameterized by a fixed set of Bézier coefficients $\alpha^{*}$. This solution is symmetrically remapped to produce a full gait cycle with alternating stance legs, yielding reference trajectories for the complete walking motion.
III-B CLF-Based Reward Design
Let $\eta(t)$ denote the output tracking error between the current state and the reference trajectory:
$\eta(t)=\begin{pmatrix}y^{d}(t)-y(q)\\ \dot{y}^{d}(t)-\dot{y}(q,\dot{q})\end{pmatrix}$, (7)
and the corresponding control Lyapunov function is
$V(\eta)=\eta^{\top}P\eta.$ (8)
The matrix $P\succ 0$ is the unique solution to the Continuous-Time Algebraic Riccati Equation (CARE) for the feedback linearized model (see below). Prop. 1 proves that this is a valid CLF for the nonlinear system in single support flat-footed walking.
Proposition 1 (CLF Existence).
Consider the fully-actuated humanoid system given by Eq. (3) and satisfying the constraint given by Eq. (4) and friction cone constraints. If $Q\succ 0$ in the CARE, and the jacobian of the virtual constraints $\frac{\partial y}{\partial q}$ are invertible, then the function $V(\eta)=\eta^{T}P\eta$ is a valid CLF for the full nonlinear dynamics during the single support phase.
Proof.
If the holonomic constraints, Eq. (4), are satisfied, and the friction cone is respected, then we can write the system dynamics without the constraint Jacobian [grizzle_3d_2010]: $D(q)\ddot{q}+H(q,\dot{q})=Bu.$ Define the virtual constraints error as $\bar{y}=y^{d}(t)-y(q)$ which we want to drive to 0. Differentiating twice yields
$\ddot{\bar{y}}=\underbrace{\frac{\partial L_{f}\bar{y}}{\partial q}\dot{q}+\frac{\partial\bar{y}}{\partial q}\left(D(q)^{-1}(-H(q,\dot{q})\right)}_{L_{f}^{2}\bar{y}(q,\dot{q})}+\underbrace{\frac{\partial\bar{y}}{\partial q}\left(D(q)^{-1}Bu\right)}_{L_{g}L_{f}\bar{y}(q,\dot{q})u}.$
From the fully actuated assumption, $\text{dim}\>q=\text{dim}\>u=\text{rank}\>B$ [grizzle_3d_2010]. Therefore $L_{g}L_{f}\bar{y}(q,\dot{q})$ is invertible under the assumptions of the proposition. We can now apply the feedback linearizing controller: $u=K(x)$, where $K(x)$ is defined as $K(x):=L_{g}L_{f}\bar{y}(q,\dot{q})^{-1}\left(-L_{f}^{2}\bar{y}(q,\dot{q})+v\right)$ where $v$ is a new, auxiliary, input. This leads to the closed loop system $\dot{\eta}=A_{\eta}\eta+B_{\eta}v$ with $A_{\eta}$ and $B_{\eta}$ defined as $A_{\eta}=\begin{bmatrix}0&I\\ 0&0\end{bmatrix}\quad B_{\eta}=\begin{bmatrix}0\\ I\end{bmatrix}$. This is a double integrator system. We solve the CARE for this system and use the resulting $P$ to synthesize an LQR gain: $K_{\eta}=R_{\eta}^{-1}B_{\eta}^{T}P$. We can then define $v=K_{\eta}\eta$ as an optimal stabilizing auxiliary controller. Denote the closed loop $\eta$ system as $A_{\eta}^{\text{cl}}=A_{\eta}+B_{\eta}K_{\eta}$. Consider the function $V(\eta)=\eta^{T}P\eta$. Using the feedback controller and differentiating yields
$\dot{V}(\eta)=\frac{\partial V}{\partial\eta}\dot{\eta}=2\eta^{T}PA_{\eta}^{\text{cl}}\eta=\eta^{T}(A_{\eta}^{\text{cl}T}P+PA_{\eta}^{\text{cl}})\eta.$ (9)
Then, using the CARE, we get that $A_{\eta}^{\text{cl}T}P+PA_{\eta}^{\text{cl}}=-Q-PB_{\eta}R_{\eta}^{-1}B_{\eta}^{T}P:=\bar{Q}$. By assumption $Q$ is positive definite and note that the other term is positive semi-definite. Therefore
$\dot{V}(\eta)\leq-\lambda_{\text{min}}(\bar{Q})\|\eta\|^{2}\leq-\frac{\lambda_{\min}(\bar{Q})}{\lambda_{\min}(P)}V(\eta)$ (10)
where $\lambda_{\min}$ denotes the minimum eigenvalue. Using $V(\eta)$ as a Lyapunov function, the set $\{u\>|\>\dot{V}\leq-\lambda V\}$ is non-empty and thus $V(\eta)$ is a CLF for the nonlinear system. ∎
Let $\sigma_{v}=\mu_{\max}(P)\,\eta_{\max}^{2}$, where $\mu_{\max}(P)$ is the maximum eigenvalue of the CLF matrix $P$ and $\eta_{\max}$ is an empirical bound on the tracking error. With $V_{t}=V\bigl(\eta(t)\bigr)$, the CLF tracking reward is then
$r_{v}$ $=w_{v}\exp\!\left(-\frac{V_{t}}{\sigma_{v}}\right).$ (CLF Tracking)
We formulate the CLF decay reward using a clip function:
$r_{\dot{v}}^{\text{clip}}$ $=-w_{\dot{v}}\;\operatorname{clip}\!\left(\tfrac{\dot{V}_{t}+\lambda V_{t}}{\sigma_{\dot{v}}},\,0,\,1\right),$ (CLF Decay)